# Analysis of Hierarchical AoII over unreliable channel: A Stochastic Hybrid System Approach

### v1

@@ Abstract
In this work, we generalize the Stochastic Hybrid Systems (SHSs) analysis of traditional AoI to the AoII metric. Hierarchical ageing processes are adopted using the continuous AoII for the first time, where two different hierarchy schemes, i.e., a hybrid of linear ageing processes with different slopes and a hybrid of linear and quadratic ageing processes, are considered. We first modify the main result in [1, Theorem 1] to provide a systematic way to analyze the continuous hierarchical AoII over unslotted real-time systems. The closed-form expressions of average hierarchical AoII are obtained based on our Theorem 1 in two typical scenarios with different channel conditions, i.e., an M/M/1/1 queue over noisy channel and two M/M/1/1 queues over collision channel. Moreover, we analyze the stability conditions for two scenarios given that the quadratic ageing process may lead to the absence of stationary solutions. Finally, we compare the average age performance between the classic AoI results and our AoII results in the M/M/1/1 queue, and the effects of different channel parameters on AoII are also evaluated.
Index Terms:
Stochastic hybrid system, continuous AoII, hierarchy scheme, stability analysis.
@@ I Introduction
In real-time status update systems, the study of information freshness over the network layer has been the subject of ongoing interest in the past decade, where a classic measure known as Age of Information (AoI) was proposed in [2, 3, 4]. Since then, extensively researches have focused on two dominant directions in this field: i). the analysis of closed-form average AoI (and its moments) over queueing and other networking systems, e.g., [5, 6, 7, 8, 9, 1, 10, 11]; ii). the optimization of AoI performance that searches for the optimal scheduling policies over different networking systems, e.g., [12, 13, 14, 15, 16, 17, 18]. These seminal works also facilitate the cross-fertilization of AoI and other related fields (e.g., topics in Information Theory), and a comprehensive survey of these works can be found in [19].
@@
However, the major drawback of AoI is that this metric only considers the time aspect of update but is content-agnostic, which may pose a challenge to identify the significance of update currently used in the system. According to the basic definition of AoI, the ageing of an update starts since it is generated and continues until it is replaced by the next successful delivered update. However, this kind of ageing measurement may not suitable for all the real-time update system. For example, if the newly generated update contains the same information as the one stored at the monitor, the ageing process of the system should be “paused” for a while since the current update is still fresh enough to be used. Therefore, a variety of modified age metrics have been proposed to address this essential problem, most of which jointly consider the content-mismatch between the transmitter and monitor by introducing suitable error functions as well as the ageing process measured by AoI and its functionals. To our best knowledge, these new metrics include Effective AoI (EAoI) [20], Age of Synchronization (AoS)[21], Query AoI (QAoI)[22], binary freshness[23], Urgency of Information (UoI)[24], and most importantly, Age of Incorrect Information [^1] There are also several variations of AoII, e.g., Age of Changed Information (AoCI)[25], Age of Incorrect Estimates(AoIE)[26], Age of Outdated Information (Ao2I)[27], Age of Processed Information (AoPI)[28]. (AoII) [29]. In fact, it is not difficult to conclude that these metrics could be classified into a more generalized form:
$x(t)=F\Big{(}f(\Delta(t)),g(X(t),\hat{X}(t))\Big{)},$ (1)
where $f:[0,\infty)\to[0,\infty)$ represents the level of dissatisfaction for freshness and $g:\Omega\times\Omega\to[0,\infty)$ defines the error metric function that penalizes the content-mismatch between the stored update and the fresh one at source. Moreover, the mapping $F:[0,\infty)\times[0,\infty)\to[0,\infty)$ describes the relationship between two metrics, e.g., the product form taken in [29]. Certainly, one can select more complex forms such as the exponents when one of the metrics plays a decisive role in achieving a particular communication goal. Hence, following [29], we call the metric in (1) generalized AoII that chooses these mappings to capture the timely utility with three degrees of freedom, which enriches the traditional AoI metric.
Recently, there have been a growing interest in studying the age performance and optimal scheduling policy using AoII, e.g., [30, 31, 32, 33, 34, 38, 35, 36, 37]. As a step in this direction, we are interested in deriving and analyzing the closed-form average age performance by exploring a more general form of AoII using (1). Our work in this paper differs from these aforementioned works in two essential ways:
@@
1) We extend the AoII metric into unslotted real-time update systems using a continuous form of (1), which, to our best of knowledge, has only been used in [37]. Furthermore, we seek for the closed-form expressions of AoII instead of an optimal policy by first introducing Stochastic Hybrid Systems (SHSs) into the analysis of content-aware ageing process, which is essentially different from the analysis in [38] of the coded status update system under different ARQ schemes.
2) We first consider hierarchical ageing processes in different system states since the level of dissatisfaction for freshness would not always grow at a constant rate in practice. On the other hand, since the age and error function are deeply coupled in (1) by the mapping $F$, the penalty of content-mismatch measured by the monitor could be severer depending on the system states (see Fig. 2), which is also reflected in our hierarchical AoII metric.
@@
$\bullet$ We give the closed-form expressions of average AoII in two typical real-time scenarios: 1) an M/M/1/1 queue over noisy channel; 2) two M/M/1/1 queues over collision channel. The system diagrams are depicted in Fig. 1a and 1b, respectively.
$\bullet$ We provide a systematic way to analyze the continuous hierarchical AoII over unslotted real-time systems using SHSs, where two different hierarchies are utilized. The first is a hybrid of linear functions with different slopes, and the second is a hybrid of linear and quadratic functions. The SHSs approach is more powerful than the traditional renewal-reward approach since the average AoII with quadratic forms cannot be easily obtained by calculating the time averages.
$\bullet$ Considering the potential instability in quadratic ageing, we give our stability conditions in Theorem 2 and 3 for two different systems to ensure the existence of stationary average AoII in terms of the second hierarchy scheme.
$\bullet$ The average age performance over the M/M/1/1 queue measured by AoII is compared with several classic results in traditional AoI in Section V-A. A few insights are obtained based on the inherent differences between AoI and AoII: i). the ceasing of age growth using AoII while the content is matched sharply cuts down the average age in the M/M/1/1 queue compared to AoI results; ii). the average age using AoII can even be smaller than a generate-at-will queue with the same Poisson service rate for some content-change probability $p$; iii). the monotonicity of the average age with the normalized utilization factor $\rho=\frac{\lambda}{\mu}$ is different between two metrics.
$\bullet$ The average AoII performance under different channel parameters in two scenarios are evaluated. In the noisy channel, the average AoII grows monotonically with $p_{e}$ regardless of utilization factors $\rho$ and content-changed probabilities $p$ and the range of $p_{e}$ is limited due to the stability issue. Meanwhile, the average AoII grows monotonically with larger $\rho$ in the collision channel, since a higher arrival rate of any source shall bring a higher collision probability where the mismatch cannot be eliminated and the age keeps growing.
$\bullet$ The computation complexity using Theorem 1 to obtain more general results of multi-access networks is discussed in Section VI, where the number of discrete states grows at least four times the node number $M$.
@@ II System Model
We first give basic assumptions for the noisy channel case with single source and the collision channel case with two sources. Then we present a brief introduction of SHSs and relevant settings for our models.
@@ II-A1 The Splitting Poisson Process
We track the average age performance of a source node observing a specific information process $X(t)$. The node can be viewed as a sensor in IoT networks or an observing unit in vehicular networks. The updates $\{X_{t}:t\in\mathbb{R}\}$ are sampled from $X(t)$ following a Poisson process with rate $\lambda$. In this work, we consider a memoryless source model, where each newly sampled update $X_{t}$ would change its content with probability $p$ and remain unchanged with probability $q=1-p$. Consequently, the sampling (arriving) process at the transmitter splits into two Poisson streams with rates $p\lambda$ and $q\lambda$.
The contender source in the collision case is denoted as $X^{c}(t)$, whose updates $\{X^{c}_{t}:t\in\mathbb{R}\}$ are sampled by a node at a Poisson rate $\lambda_{c}$. We assume that the content of $X_{t}$ is independent of the content of $X^{c}_{t}$, therefore, the age behavior of $X(t)$ is only related to whether or not a collision arises.
@@ II-A2 Transmission Through Unreliable Channels
Since time is unslotted, a transmission immediately starts when a newly sampled update arrives. To avoid a combinatorial explosion of the state space, we first assume the transmission times of the updates are modeled as independent exponential $\mu$ and $\mu_{c}$ random variables for updates $X_{t}$ and $X^{c}_{t}$, respectively. Furthermore, the transmission of updates is preemptive considering that the content-changed updates should be delivered as soon as possible, and the preempted update would be discarded. Hence, the transmission at each node follows M/M/1/1.
@@
The channel is unreliable in both cases. In Fig. 1a, an update may be decoded incorrectly by the monitor due to channel noise. We denote $p_{c}$ as the probability of an update being correctly decoded, and an update that is not decoded correctly is lost with probability $p_{e}=1-p_{c}$. Note that only a content-changed update $X_{t}$ being correctly decoded can reduce the age at monitor.
@@
In Fig. 1b, the channel is assumed to be noise-free yet suffering from collision. A collision period begins if a newly sampled update $X^{c}_{t}$ / $X_{t}$ arrives while there is an update from $X_{t}$ / $X^{c}_{t}$ currently being transmitted. The collision disrupts transmission in the way that neither of updates can be successfully decoded by the monitor. The update finishing transmission first would be discarded, whereas the remaining update would keep transmitting. The transmission time of the remaining update is still an exponential random variable due to the memoryless property.
II-A3 End-to-End Interaction and Age Measurements
@@
There are two common assumptions shared in real-time status-updating networks: i). the amounts of communication overheads including controlling signals require negligible transmission time; ii). these overheads are transmitted via a dedicated control channel to ensure the signaling free from errors[39]. Based on that, we can further assume that the indicator of content changes at source is immediately known at the monitor, which plays a key role in characterizing the evolution of our hierarchical AoII. To track the age of $X(t)$, we specify the generalized form (1) in the following:
$x(t)=f(t-u_{t})I\{X_{t}\neq\hat{X}_{t}\},\ t\in\mathbb{R},$ (2)
where $\hat{X}_{t}$ is the last successfully decoded update stored at the monitor, and $u_{t}$ denotes the sampling time of $\hat{X}_{t}$. The indicator function $I$ takes values from 0 or 1 if the content of update being transmitted is the same or not as $\hat{X}_{t}$. Furthermore, we assume that $I$ equal to 0 if there is no update on transmission and the content of source is not changed. The age function $f$ can be linear or quadratic depending on the SHSs state.
@@
II-B A Brief Introduction to SHSs
@@
$\mathbf{\dot{x}}=\frac{\partial\mathbf{x}}{\partial t}=f^{\prime}(\mathbf{q},\mathbf{x},t)+g^{\prime}(\mathbf{q},\mathbf{x},t)\frac{\partial\mathbf{z}}{\partial t},$ (3)
where the system states are partitioned into the discrete state $\mathbf{q}\in\mathbb{Q}$ consistent with the ordinary jump process, and a $\mathit{k}$-dimensional stochastic process $\mathbf{x}$ with piecewise continuous sample paths called the continuous state. [^2] For notation simplicity, we here omit the parameter $t$ associated with the joint system state $(\mathbf{q}(t),\mathbf{x}(t))$ at time $t$. Given a discrete state space $\mathbb{Q}$ and a $\mathit{k}$-dimensional vector $\mathbf{z}$ of independent Brownian motion processes, the continuous state $\mathbf{x}$ evolves with two state-dependent functions $f^{\prime}$ and $g^{\prime}$. Moreover, there is a family of discrete transition maps that tracks possible changes of the joint system state $(\mathbf{q},\mathbf{x})$, i.e.,
@@
Based on the general settings of SHSs above, the continuous state can be defined as the age processes for our AoII models, i.e., $\mathbf{x}(t)=[x_{1}(t),x_{2}(t)]$. We denote $x_{2}(t)$ as the AoII of $X_{t}$ at the monitor, and $x_{1}(t)$ is what the AoII at the monitor would become if an update in service were to complete transmission at time $t$. Since the AoII evolves deterministically at each state $\mathbf{q}$, we have $g^{\prime}(\mathbf{q},\mathbf{x},t)=0$ and $\mathbf{z}=0$. Moreover, the transition rates $\lambda_{l}$ of SHSs are all constant Poisson rates defined above according to the M/M/1/1 queue discipline.
Above all, the hierarchy of AoII of $X_{t}$ can be demonstrated by a piecewise age function $f$, whose derivative is denoted as $f^{\prime}$ in (4). The discrete state $\mathbf{q}$ of SHSs is the system state. Both would be later specified for each case.
$t$$x_{2}(t)$$1_{m}$$1_{m}$$1_{u}$$0_{e}$$0_{c}$$t_{0_{c}}$
$1_{m}$$1_{m}$$0_{c}$$1_{u}$$0_{e}$$t$$x_{2}(t)$$t_{0_{c}}$
Figure 2: The evolution of AoII $x_{2}(t)$ at monitor in the noisy channel case. (a) the linear age penalty with changes of slope in states $1_{u}$ and $0_{e}$; (b) the quadratic age penalty in states $1_{u}$ and $0_{e}$.

### v2

# Analysis of Hierarchical AoII over Unreliable Channels: A Stochastic Hybrid System Approach

@@ Abstract
In this work, we generalize the Stochastic Hybrid Systems (SHSs) analysis of Age of information (AoI) to the Age of Incorrect Information (AoII) metric. Hierarchical ageing processes are adopted using the continuous AoII for the first time. Two different hierarchy schemes are considered: 1). A hierarchy of zero and linear ageing processes with different slopes; 2). A hierarchy of zero, linear and exponential ageing processes. We first modify the main result in [1, Theorem 1] to provide a systematic way to analyze the continuous hierarchical AoII over unslotted real-time systems. The closed-form expressions of average hierarchical AoII are obtained in two typical scenarios with different channel conditions, i.e., an M/M/1/1 queue over noisy channels and two M/M/1/1 queues over collision channels. Moreover, under each scenario, we analyze the stability issue regarding positive recurrence and provide the stability conditions that ensure a steady-state average AoII. Finally, we compare the closed-form results between average AoI and AoII in the M/M/1/1 queue. The effects of different channel parameters on the average hierarchical AoII are also evaluated.
@@ I Introduction
In real-time status update systems, the study of information freshness over the network layer has been the subject of ongoing interest in the past decade. A classic measure, Age of Information (AoI), was first introduced to measure information freshness. The AoI $\Delta(t)$ measured by the monitor at time $t$ is defined by $\Delta(t)=t-U(t)$, which is the elapsed time since the generation time $U(t)$ of the last received update. Since then, extensively researches have focused on two dominant directions in the AoI field: i). The analysis of closed-form average AoI (and its moments) over queueing and other networking systems, e.g., [2, 3, 4, 5, 6, 1, 7, 8]; ii). The optimization of AoI performance that searches for the optimal scheduling policies over different networking systems, e.g., [9, 10, 11, 12, 13, 14, 15]. A comprehensive survey of these works can be found in [16].
@@
However, the major drawback of AoI is that this metric only considers the time aspect of update but is content-agnostic. The AoI process keeps growing unless a new update is received. However, such an ageing process may not suitable for all the real-time update systems. For example, if the newly-generated source update contains the same information as the one stored at the monitor, the ageing process of the system should be “paused” for a while since the current update is still fresh enough to be used. Therefore, a variety of modified metrics have been proposed to address this content-agnostic issue. To our best knowledge, these new metrics include Effective AoI (EAoI) [17], Age of Synchronization (AoS)[18], Query AoI (QAoI)[19], binary freshness[20], Urgency of Information (UoI)[21], Age of Incorrect Information (AoII) [22], etc. [^1] There are also several variations of AoII, e.g., Age of Changed Information (AoCI)[23], Age of Incorrect Estimates(AoIE)[24], Age of Processed Information (AoPI)[25]. These metrics not only include the age perspective but also consider the content-mismatch between the updates at the source and monitor. In this work, we adopt the most representative form among them, i.e., AoII, and extend the SHSs analysis of AoI [6] into the AoII metric. First, we give the definition of AoII first proposed in [22], i.e.,
$x(t)=f(t)\times g(X(t),\hat{X}(t)),$ (1)
where function $f:[0,\infty)\to[0,\infty)$ is an increasing age penalty function paid for being unaware of the correct status of the source process. Function $g:\Omega\times\Omega\to[0,\infty)$ is an information penalty function that reflects the content-mismatch between the current estimate $\hat{X}(t)$ at the monitor and the actual content $X(t)$ of the process. Hence, AoII is identified as a content-aware age metric and also a promising alternative measurement of information freshness compared to AoI. There has been a growing interest in studying the average AoII performance and optimal scheduling policy, e.g., [26, 27, 28, 29, 30, 31, 32, 33, 34]. However, our work differs from them in two essential ways:
@@
1) We extend the AoII metric into unslotted real-time update systems using a continuous form of (1), which, to our best of knowledge, has only been used in [33]. Furthermore, we pursue the closed-form expressions of AoII instead of an optimal policy by first introducing Stochastic Hybrid Systems (SHSs) into the analysis of content-aware AoII process.
2) Since the level of age dissatisfaction measured in (1) is determined by different choices for functions $f$ and $g$, it will not always grow at a constant rate (as AoI) in practice. Hence, we first consider hierarchical AoII processes, defined in our Definition 1. That is to say, in this work, the ageing process using AoII depends on the system state, which can have a zero, or linear, or exponential growth, thus hierarchy, at different system states.
@@
$\bullet$ We give the closed-form expressions of average hierarchical AoII in two typical real-time scenarios: 1). A source delivers updates to a monitor over a noisy channel, where the arrival and transmission of source updates are formulated as an M/M/1/1 queue; 2). A source of interest and a contender source both deliver updates to a monitor over a collision channel, where the arrival and transmission of updates from each source are formulated as an M/M/1/1 queue, respectively. The system diagrams are depicted in Fig. 1a and 1b.
$\bullet$ We provide a systematic way to analyze the continuous hierarchical AoII over unslotted real-time systems using SHSs, where two kinds of hierarchy schemes are discussed. The first is a hierarchy of zero and linear growth of AoII processes, where the linear growth of AoII has different slopes at different states. The second is a hierarchy of zero, linear, and exponential growth of AoII processes.
$\bullet$ Following [6] and [35], we discuss the stability issue of hierarchical AoII under SHSs. We first prove a weak-sense stability, i.e., positive recurrence, for the AoII SHSs in Lemma 2. Then, we give the steady-state stability conditions in Theorem 2 and 3 for the two different scenarios in this work, receptively, which ensure the existence of average AoII.
$\bullet$ In Section V-A, several classical results of average AoI under the M/M/1/1 queue model are compared with the average AoII in this work. A few insights are obtained based on the inherent differences between AoI and AoII: i). The ceasing of ageing using AoII sharply cuts down the average age in the M/M/1/1 queue compared to the results using AoI; ii). The average age using AoII can even be smaller than a generate-at-will queue with the same Poisson service rate $\mu$ for some content-change probability $p$; iii). The monotonicity of the average age with the normalized utilization factor $\rho=\frac{\lambda}{\mu}$ is different between these two age metrics.
$\bullet$ The average AoII performance under different channel parameters is evaluated. In the noisy channel case, the average AoII grows monotonically with $p_{e}$ regardless of the utilization factor $\rho$ and content-changed probabilities $p$. In the collision channel case, the average AoII grows monotonically with a larger $\rho$ since a higher arrival rate of updates from each source gives a higher collision probability. The content mismatch cannot be eliminated as soon as possible due to collisions, which results in a larger average AoII.
$\bullet$ The SHSs analysis can be generalized to $M\geqslant 3$ sources. The computation complexity using Theorem 1 to obtain more general results of multi-access networks is discussed in Section VI, where the number of discrete states grows at least four times the number of nodes $M$.
@@ II System Model
We first give our basic assumptions for the noisy channel case with single source and the collision channel case with two sources. Then we present the basic settings and relevant definitions for the AoII SHSs.
@@ II-A1 The Splitting Poisson Process
We track the average AoII performance of a source node observing a specific information process $X(t)$. The node can be viewed as a sensor in IoT networks or an observing unit in vehicular networks. The updates $\{X_{t}:t\in\mathbb{R}\}$ are sampled from $X(t)$ following a Poisson process with rate $\lambda$. In this work, we consider a memoryless source model, where each newly-sampled update $X_{t}$ changes its content with probability $p$ and remain unchanged with probability q=1pq=1-p. The sampling (arriving) process at the transmitter therefore splits into two Poisson streams with rates $p\lambda$ and $q\lambda$. Moreover, following the discussion in [31, 34, 36], we assume that the state space of the content of $X_{t}$ is very large (possibly infinite) and unknown, where the likelihood of content returning to a previous state is negligible if it has changed its state.
The contender source in the collision case is denoted as $X^{c}(t)$, whose updates $\{X^{c}_{t}:t\in\mathbb{R}\}$ are sampled by a node at a Poisson rate $\lambda_{c}$. We assume that the content of $X_{t}$ is independent of the content of $X^{c}_{t}$. Therefore, the AoII process of $X(t)$ measured at the monitor is only affected by $X^{c}_{t}$ in a sense that whether or not a collision arises.
@@ II-A2 Transmission Through Unreliable Channels
Since time is unslotted, a transmission immediately starts when a new update arrives. We assume the transmission times of the updates are modeled as independent exponential $\mu$ and $\mu_{c}$ random variables for updates $X_{t}$ and $X^{c}_{t}$, respectively. Furthermore, the transmission of updates is preemptive considering that the content-changed updates should be delivered as soon as possible. Hence, we assume that the transmission of updates from $X(t)$ and $X^{c}(t)$ both follows M/M/1/1, where the preempted updates are discarded.
@@
The channel is unreliable in both cases. In Fig. 1a, an update may be decoded incorrectly by the monitor due to channel noise. We denote $p_{c}$ as the probability of an update being correctly decoded, and an update that is not decoded correctly is lost with probability $p_{e}=1-p_{c}$. Note that only an update $X_{t}$ being correctly decoded can reduce the AoII at the monitor.
@@
In Fig. 1b, the channel is assumed to be noise-free yet suffering collisions. A collision period begins if a newly sampled update $X^{c}_{t}$ / $X_{t}$ arrives while there is an update from $X_{t}$ / $X^{c}_{t}$ currently being transmitted. The collision period ends if one of these updates finishes transmission. Hence, the collision disrupts transmission in the way that the update finishing transmission first cannot be successfully decoded and being discarded, whereas the remaining update would keep transmitting. The transmission time of the remaining is still exponential due to the memoryless property. This consideration leads to two different cases that the remaining update can or cannot be successfully decoded, both of which are addressed in Section IV.
II-A3 End-to-End Interaction
@@
There are two common assumptions shared in real-time status-updating networks: i). The amounts of communication overheads including feedback signals require negligible transmission time; ii). These overheads are transmitted via a dedicated control channel to ensure the signaling free from errors [^2] To ensure reliability in such a control channel, a reliable protocol can be employed[37].. Since the state space of the content of $X_{t}$ is assumed to be unknown, our source model in fact cannot require the knowledge of source content, but rather the knowledge of a one-bit indicator of content changing at source. This can be implemented under locally-deployed processors, e.g., pre-trained neural networks, at the source side that output such an indicator. Furthermore, we assume that the indicator of content-change is instantaneously known at the monitor based on i) and ii), which results in changes of system states.
Remark 1.
Inspired by [28], we discuss a potential way to implement such an indicator, where some locally-deployed processors, e.g., pre-trained neural networks, at the source side are used to output the indicator. Define $I_{s}=\mathbb{I}\{c(X_{t_{1}},X_{t_{0}})\geqslant d\}$, where $\mathbb{I}\{\cdot\}$ is an indicator function taking values from 0 and 1. The function $c(.,.)$ is some cost function that measures the content mismatch between two updates, and $d$ is some predefined threshold. Note that $c(.,.)$ can be designated or fitted by the pre-trained neural network. The function $f$ determines whether a content change exists between $X_{t_{1}}$ and the last sample $X_{t_{0}}$. When $c(X_{t_{1}},X_{t_{0}})\geqslant d$, $I_{s}=1$ triggering a content-change indicator to the monitor; otherwise, $I_{s}=0$. The threshold parameter $d$ controls the system’s tolerance to content mismatches, which is adjustable for different sources.
@@
II-B Basic Settings of AoII SHSs
@@
$\mathbf{\dot{x}}=\frac{\partial\mathbf{x}}{\partial t}=f^{\prime}(\mathbf{q},\mathbf{x},t)+g^{\prime}(\mathbf{q},\mathbf{x},t)\frac{\partial\mathbf{z}}{\partial t}.$ (2)
The system state $(\mathbf{q}(t),\mathbf{x}(t))$ is partitioned into the discrete state $\mathbf{q}\in\mathbb{Q}$ consistent with the ordinary jump process, and a $\mathit{k}$-dimensional stochastic process $\mathbf{x}$ with piecewise continuous sample paths called the continuous state. Given a discrete state space $\mathbb{Q}$ and a $\mathit{k}$-dimensional vector $\mathbf{z}$ of independent Brownian motion processes, the continuous state $\mathbf{x}$ evolves with two state-dependent functions $f^{\prime}$ and $g^{\prime}$. Moreover, there is a family of discrete transition maps that tracks possible changes of the joint system state $(\mathbf{q},\mathbf{x})$,
@@
Based on the general settings of SHSs above, the continuous state can be defined as the AoII processes of interest in our models where $\mathbf{x}(t)=[x_{1}(t),x_{2}(t)]$. We denote $x_{2}(t)$ as the AoII of $X_{t}$ at the monitor, and $x_{1}(t)$ is what the AoII at the monitor would become if an update in service were to complete transmission at time $t$. Since the AoII evolves deterministically at each state $\mathbf{q}$, we have $g^{\prime}(\mathbf{q},\mathbf{x},t)=0$ and $\mathbf{z}=0$. Moreover, the transition rates $\lambda_{l}$ of the AoII SHSs are all constant Poisson rates according to the system model.
Above all, the hierarchy of AoII of $X_{t}$ is characterized by the state-dependent function $f^{\prime}(\mathbf{q},\mathbf{x},t)$, which is called the growth rate function in our paper. Since we adopt SHSs to analyze our model using the continuous AoII (1), the growth rate function $f^{\prime}(\mathbf{q},\mathbf{x},t)$ can be interpreted as the derivative of AoII at different discrete state $\mathbf{q}$. For example, if we have the same content between the source and monitor at some state $\mathbf{q}_{1}$, then $f^{\prime}(\mathbf{q}_{1},\mathbf{x},t)=0$. However, if we use a constant growth rate at some state $\mathbf{q}_{2}$ when there is a content-mismatch, then $f^{\prime}(\mathbf{q}_{2},\mathbf{x},t)=m$ where $m$ is a positive constant. That is, the ageing process at $\mathbf{q}_{2}$ grows like AoI with slope $m$. Moreover, if we use a linear growth rate [^3] The terminology follows from [38, Lemma 1] in $\mathbf{x}$ at some state $\mathbf{q_{3}}$ when the content-mismatch is considered to be severe, then $f^{\prime}(\mathbf{q}_{3},\mathbf{x},t)=k\mathbf{x}$. Then, the ageing process grows exponentially at $\mathbf{q_{3}}$. To that end, we can define the concept hierarchical AoII used in this paper.
Definition 1 (Hierarchical AoII).
Under SHSs model, the hierarchical AoII using the continuous form (1) has different growth rate functions $f^{\prime}(\mathbf{q},\mathbf{x},t)$ at different system states.
In order to reveal closed-form results of the average hierarchical AoII, we adopt an indicator function $I=\mathbb{I}\{X_{t}\neq\hat{X}_{t}\}$ to specify the information penalty function $g$ in (1), which is a common consideration in the AoII literature. The indicator function $I$ takes values from 0 or 1 if the content of update being transmitted is the same or not as the estimate $\hat{X}_{t}$, i.e., the last successfully decoded update, stored at the monitor. Hence, such an indicator is state-dependent in AoII SHSs, i.e., $I=I(\mathbf{q})$. We also note that, for the memoryless source with zero returning probability, using an indicator of the content-change between two successive source updates is the same as using $I(\mathbf{q})$. Moreover, based on Definition 1, the age penalty function $f$ that adopts to characterize hierarchical AoII can have a linear or exponential growth depending on the choice of $f^{\prime}(\mathbf{q},\mathbf{x},t)$ for different $\mathbf{q}$. So, we can rewrite (1) as
$x(\mathbf{q},\mathbf{x},t)=f(\mathbf{q},\mathbf{x},t)I(\mathbf{q}),\ t\in\mathbb{R}.$ (4)
Above all, since we leverage SHSs to formulate our models, in the following, we characterize the specified hierarchical AoII processes $x(\mathbf{q},\mathbf{x},t)$ by adopting different hierarchical schemes of $f^{\prime}(\mathbf{q},\mathbf{x},t)$ under the considered scenarios.
II-C Stability Analysis of SHSs - Fundamentals
In studying SHSs, researchers use Lagrange stability to depict systems that are ultimately uniformly bounded in the $n$-th mean ($n\geqslant 1$) of $\mathbf{x}(t)$[39, 40, 35]. However, as suggested in [39, Chapter 3], for systems that only focus on the time and ensemble average of $\mathbf{x}(t)$, it is more appropriate to analyze a weak-sense stability, i.e., positive recurrence, which encloses the case $n=1$ in Lagrange stability.
Definition 2 (Recurrence and Positive Recurrence[39]).
For any set $\mathbb{V}=\mathbb{Q}\times D\subset\mathbb{Q}\times\mathbb{R}^{n}$ and an initial condition $(\mathbf{q}_{0},\mathbf{x}_{0})\in\mathbb{V}^{c}$, the complement of $\mathbb{V}$, the recurrent time is the time interval that $(\mathbf{q},\mathbf{x})$ at time $t$ returns to $\mathbb{V}$ from the initial state, i.e.,
$\sigma_{\mathbb{V}}=\inf\{t:(\mathbf{q}_{0},\mathbf{x}_{0})\in\mathbb{V}^{c},(\mathbf{q},\mathbf{x})\in\mathbb{V}\},$ (5)
A regular SHS process is recurrent with respect to $\mathbb{V}$ if
$\mathrm{Pr}[\sigma_{\mathbb{V}}<\infty]=1,\ \forall\ (\mathbf{q}_{0},\mathbf{x}_{0})\in\mathbb{V}^{c}.$ (6)
In addition, a recurrent process with finite mean recurrence time for some $\mathbb{V}$ is positive recurrent with respect to $\mathbb{V}$.
A regular SHS process is an SHS that always has a unique global solution for any initial condition. We assume that the regularity holds for the AoII SHSs from now on, which is verified later in the proof in Appendix B.
The Lagrange stability of SHSs is discussed in depth by Maatouk et al. in [35]. They strictly prove the positive recurrence of the proposed age-aware SHSs and also prove the Lagrange stability for the higher-order age moments. Following them, we give our stability analysis of our AoII SHSs first from the perspective of positive recurrence since we are only interested in the average AoII.
In addition, we discuss the necessary conditions that are determining factors to ensure the existence of steady-state (or ergodic) average AoII after we obtain the positive recurrence. The analysis of such conditions was first carried out by Yates in [6, Section V]. Following this, we then analyze the steady-state stability conditions of the AoII SHSs based on ergodicity.
[…]
Figure 2: Sample paths of AoII $x_{2}(t)$ at monitor in the noisy channel case under: (a) hierarchy scheme 1; (b) hierarchy scheme 2. $t_{0_{c}}$ is the time instance that the content is matched.